1. How many moles of ions are in the following solutions: 32.5mg potassium sulfate in water; 1 mol MgBr2 in C6H6?

2. Given the following reaction and solubility data, how many moles of total ions are in solution once the reaction is complete starting with 25g of each reactant in 1L?
Na3PO4 + Ca(NO3)2 --> NaNO3 + Ca3(PO4)2

COMPOUND SOLUBILITY
SODIUM PHOSPHATE 94.6g/100mL
CALCIUM NITRATE 120g/100mL
SODIUM NITRATE 91.2g/100 mL
CALCIUM PHOSPHATE 2mg/100mL
Answer #1

1. Mass of K2SO4 = 32.5 mg = 32.5 *10-3 g = 0.0325 g

Molar mass of K2SO4 = 174.259 g/mol

Moles of K2SO4 = mass/ Molar mass = 0.0325 g / ( 174.259 g/mol) = 1.865 *10-4 mol

Thus, moles of K+ = 2* 1.865 *10-4 mol = 3.73 *10-4 mol

moles of SO42- = 1.865 *10-4 mol

Hence the total moles of ions produced = (1.865 *10-4 mol )+ (3.73 *10-4 mol) = 5.60 *10-4 mol

(b) MgBr2 is a polar but C6H6 is non- polar

SO both are insoluble

Hence, there is no ion

2.

Reaction is:

Na3PO4 + Ca(NO3)2 --> NaNO3 + Ca3(PO4)2
Mass of Na3PO4 = 25 g

Molar mass of Na3PO4 = 163.94 g/mol

Moles of Na3PO4 = 25/ 163.94 = 0.152 mol

Mass of Ca(NO3)2 = 25 g

Molar mass of Ca(NO3)2 = 164.088 g/mol

Moles of Ca(NO3)2 = 25 /164.088 g/mol = 0.152 mol

Thus,Moles of NaNO3 = moles of Ca3(PO4)2 = 0.152 mol

From solubility , Ca3(PO4)2 is insoluble. so, didn;t ionize

Only , NaNO3 will ionize into Na+ and NO3- = 2 ions

hence total moles of ions formed = 2* 0.152 = 0.304 moles